a + 9b=4
a + 4b= -1
solve by the elimination method

Found 2 solutions by Alan3354, checkley79:
Answer by Alan3354(69443)

a + 9b=4
a + 4b= -1
---------------- Subtract

Answer by checkley79(3341)

a + 9b=4
a + 4b= -1 SUBTRACT.
-----------------------------
5b=5
b=5/5
b=1 ans.
a+9*1=4
a+9=4
a=4-9
a=-5 ans.
proof"
-5+4*1=-1
-5+4=-1
-1=-1